# Towards Activated Muscle Group Estimation in the Wild

### v4

@@ 4.7. Analysis of the Ablation Studies
Module ablation. The ablation study of MCT, MCTKD, and MCTF, is shown in Table 3, where we deliver the results for w/o MCT, w/o MCTKD, w/o MCTF, and w/ all. When we compare the results between w/o MCT and w/ all, we find that using MCT to enlarge the attributes prediction space can contribute performance improvements by $2.1\%$, $2.0\%$, $2.1\%$, $1.9\%$, $2.2\%$, and $2.0\%$ in terms of known val, new val, mean val, known test, new test, and mean test. When comparing the results between w/o MCTKD and w/ all, we observe that leveraging MCTKD to achieve multi-stage information exchange from the video model and skeleton model can harvest performance improvements by $2.1\%$, $2.0\%$, $2.1\%$, $1.9\%$, $2.2\%$, and $2.0\%$ in terms of the six aforementioned evaluations. When we compare the results between w/o MCTF and w/ all, we find that using MCTF to achieve the fusion between the information derived from the classification MCT and knowledge distillation MCT can bring performance improvements of $2.4\%$, $1.7\%$, $2.1\%$, $2.2\%$, $2.1\%$, and $2.1\%$ in terms of the six aforementioned evaluations.
@@
Figure 5. Qualitative results for the MViTv2-S (Li et al., 2022) and TransM3E. GradCam (Selvaraju et al., 2017) visualization is given. The ground truth is shown on the left, and the prediction and gradients of the MViTv2-S (Li et al., 2022), and our approach are shown in the middle and on the right.
MCTKD ablation. We evaluate the effects of varying the location of knowledge distillation application and the knowledge distillation on a single distillation token (KD) and MCT (MCTKD), where they are named differently, i.e., KD/MCTKD at the final layer (FL-KD/MCTKD), KD/MCTKD after token size reduction (SP-KD/MCTKD), or KD/MCTKD after each MViTv2 block (DE-KD/MCTKD), in Table 4. Considering different knowledge distillation localizations, SP-KD and SP-MCTKD achieve the best performances for KD and MCTKD individually, demonstrating their superiority of using sparse knowledge distillation settings after the reduction of the feature map size. When we compare sparse knowledge distillation with dense knowledge distillation, SP-MCTKD achieves performance improvement by $2.7\%$, $0.8\%$, $1.8\%$, $2.6\%$, $0.8\%$, and $1.7\%$ in terms of known val, new val, mean val, known test, new test, and mean test. When we compare sparse knowledge distillation with final layer knowledge distillation, SP-MCTKD achieves performance improvements by $2.7\%$, $1.1\%$, $1.9\%$, $2.3\%$, $1.4\%$, and $1.8\%$ in terms of the aforementioned six evaluations.
Using MCTKD in sparse locations—specifically after each feature map size reduction—yields the best results on the MuscleMap benchmark. This method enhances the aggregation of AMGE cues across modalities after pooling, facilitating more effective knowledge distillation. SP-MCTKD surpasses SP-KD by utilizing multi-classification tokens to expand the prediction space for attributes during training, thereby capturing and transferring more crucial AMGE cues from various modalities. Consequently, SP-MCTKD, which excels in performance, is chosen as the foundation for our model, facilitating knowledge exchange between skeleton and video models. SP-MCTKD achieves the best performance and is selected.
MCTF ablation. The ablations on MCTF for TransM3E are presented in Table 5, where our approach is compared with existing fusion approaches, e.g., Sum, Multiplication, SelfAttention, and CrossAttention. MCTF shows the best performance with $81.0\%$ and $81.0\%$ on mean val and mean test. Compared with CrossAttention, our MCTF achieves performance improvement by $1.7\%$ and $1.7\%$ in terms of the mean val and mean test. The superiority of MCTF compared to other approaches, especially on generalizability, depends on using attention from a more diverse perspective, which benefits the capability of integration considering different focus formats.
4.8. Comparison with Conventional Multi-modality Fusion Approaches
Table 6 presents the comparison between TransM3E and existing multi-modality fusion approaches, i.e., LateFuionSum, LateFusionConcat, and LateFusionMul. We compare our proposed method to these conventional multi-modal fusion approaches to illustrate that the performance improvement of our approach is not solely delivered by using the feature fusion between the skeleton modality and the RGB video modality. Compared with the best performing baseline LateFusionConcat, our approach achieves a performance improvement by $14.3\%$, $3.3\%$, $8.8\%$, $14.5\%$, $3.0\%$, and $8.7\%$ in terms of known val, new val, mean val, known test, new test, and mean test.
@@ 5. Conclusion
In this paper, we open the vistas of video-based activated muscle group estimation in the wild. We contribute the first large-scale video-based activated muscle group estimation dataset considering in-the-wild video and build up MuscleMap benchmark for the AMGE by using statistic baselines and existing video-based approaches including both video-based and skeleton-based methods. We take additional consideration regarding the AMGE generalizability. We propose TransM3E with multi-classification token distillation and fusion in a cross-modality manner to enhance the generalization to new activity types. TransM3E sets the state-of-the-art on the proposed MuscleMap benchmark and it delivers promising generalizability towards unseen activities in terms of the AMGE in-the-wild task. In the future, the utilization of large language models (LLMs) for the purpose of video-based active muscle group estimation is anticipated to enhance the extensibility of the approach with respect to estimating muscle groups during novel physical activities.
@@ Appendix A Society Impact and Limitations
In our work, a new dataset targeting at the AMGE is collected based on YouTube videos, termed as MuscleMap135. We build up the MuscleMap benchmark for the AMGE by using statistic baselines and existing video-based approaches including both video-based and skeleton-based methods, while the three aforementioned datasets are all considered. Through the experiments we find that the generalizability targeting AMGE on new activities is not satisfied for the existing activity recognition approaches. In order to tackle this issue, we propose a new cross-modality knowledge distillation approach named TransM3E while using MViTv2-S (Li et al., 2022) as its basic backbone. The proposed approach alleviates the generalization problem in a certain degree, however, there is still a large space for further improvement and future research. The AMGE performance gap between the known activities and new activities illustrates that our model has potential to give offensive predictions, misclassification, and biased content which may cause false predictions resulting in a negative social impact. The dataset and code will be released publicly.
@@ Appendix B More details of the Dataset
The muscle regions where the number of sources are bigger than the threshold are chosen as activated muscle region. We can see that no obvious deviation could be found in the AMGE annotation. We annotate the commonly leveraged human body muscles in daily life into $20$ muscle regions according to the suggestion of the experts, i.e., neck and head region, chest region, shoulder region, biceps region, triceps region, forearms region, upper back region, latissimus region, obliques region, upper abdominis region, lower abdominis region, lower back region, hamstring region, quadriceps region, calves region, inner thigh region, outer thigh region, gluteus region, feet ankles region, and wrists region. We rearrange occipitofrontalis, temporoparientalis, levator labii superioris, masticatorii, sternocleidomastoideus as neck and head muscle region; pectoralis major as chest region; deltoideus as shoulder region; biceps brachii as biceps region; triceps brachii as triceps region; flexor carpi radialis, palmaris longus, abductor pollicis longus as forearm region; trapezius as upper back region; latissimus dorsi as latissimus region; external oblique, serratus anterior as obliques region; rectus abdominis, quadratus lumborum as upper abdominis region; transversus abdominis, pyramidalis as lower abdominis region; erector spinae as lower back region; biceps femoris, semimembranosus, semitendinosus as hamstring region; rectus femoris, vastus medialis as quadriceps region; gastrocnemius, soleus as calves region; adductor longus, sartorius, gracilis as inner thigh region; iliotibial tract as outer thigh region; gluteus maximus as gluteus region; peroneus longus and brevis, extensor digitorum longus, flexor hallucis longus, flexor digitorum longus, peroneus tertius, tibialis posterior as feet ankles region; extensor pollicis, 1st dorsal interosseous, pronator quadratus as wrists region.
@@ Appendix E More Details of the MCTKD
Since we introduced the ablation regarding MCTKD in our main paper with experimental results, only more details regarding the KD format and position will be introduced in this section. In order to make it clearer for understanding, we illustrate more details regarding the KD/MCTKD position in Figure 6 to give a detailed clarification. For the MCTKD-related approaches, we use the MCTKD as depicted by (d), where the KD is executed between the knowledge receiver MCTs of the main modality and the sender MCTs of the auxiliary modality. For all the other basic KD-based approaches, we use the format as depicted by (c), where the KD is executed between the MCTs of the main modality and the MCTs of the auxiliary modality, regarded as conventional KD. All the experiments are executed with MCTs while without MCTF aggregation. We simply average the MCTs for all the experiments in this ablation. Regarding the sparse format as depicted in (a), the knowledge of the auxiliary modality is only transferred after the size reduction of the pooling layer denoted as DownSampling (DS) in Figure 6 and after the final layer. Only SparseMCTKD and DenseMCTKD are depicted since the SparseKD and DenseKD use the same position settings. SparseKD/MCTKD aims at reducing the KD/MCTKD calculation by selecting the most important intermediate layers to transfer the knowledge. After each pooling layer which has size reduction, the informative cues will be highlighted, which makes the corresponding changes of the tokens from auxiliary modality necessary to be integrated through KD/MCTKD. We choose the position after the pooling with size reduction to do the KD/MCTKD on the intermediate layer. DenseKD/MCTKD is designed to transfer the knowledge directly after each transformer block in order to leverage the knowledge from the other modality thoroughly. We make use of both KD positions to conduct comparison and select the most appropriate method to build the MCTKD in our final model.

### v5

@@ 5.3. Analysis of the Ablation Studies
Module ablation. The ablation study of MCT, MCTKD, and MCTF, is shown in Table 3, where we deliver the results for w/o MCT, w/o MCTKD, w/o MCTF, and w/ all. When we compare the results between w/o MCT and w/ all, we find that using MCT to enlarge the attributes prediction space can contribute performance improvements by $2.1\%$, $2.0\%$, $2.1\%$, $1.9\%$, $2.2\%$, and $2.0\%$ in terms of known val, new val, mean val, known test, new test, and mean test. When comparing the results between w/o MCTKD and w/ all, we observe that leveraging MCTKD shows more benefits. When we compare the results between w/o MCTF and w/ all, we find that using MCTF to achieve the fusion between the information derived from the classification MCT and knowledge distillation MCT can bring performance improvements of $2.4\%$, $1.7\%$, $2.1\%$, $2.2\%$, $2.1\%$, and $2.1\%$ in terms of the six aforementioned evaluations.
@@
Figure 5. Qualitative results for the MViTv2-S (Li et al., 2022) and TransM3E. GradCam (Selvaraju et al., 2017) visualization is given.
MCTKD ablation. We evaluate the effects of varying knowledge distillation location and the knowledge distillation on a single distillation token (KD) and MCT (MCTKD), where they are named differently, i.e., KD/MCTKD at the final layer (FL-KD/MCTKD), KD/MCTKD after token size reduction (SP-KD/MCTKD), or KD/MCTKD after each MViTv2 block (DE-KD/MCTKD), in Table 4. SP-KD and SP-MCTKD achieve the best performances for KD and MCTKD individually, demonstrating their superiority of using sparse knowledge distillation settings after the reduction of the feature map size, and SP-MCTKD outperforms SP-KD. Using MCTKD in sparse locations—specifically after each feature map size reduction—yields the best results on the MuscleMap benchmark. This enhances the aggregation of AMGE cues across modalities after pooling, facilitating effective knowledge distillation. SP-MCTKD achieves the best performance and is selected.
MCTF ablation. The ablations on MCTF for TransM3E are presented in Table 5, where our approach is compared with existing fusion approaches, e.g., Sum, Multiplication, SelfAttention, and CrossAttention. MCTF shows the best performance with $81.0\%$ and $81.0\%$ on mean val and mean test. The superiority of MCTF compared to other approaches, especially on generalizability, depends on using attention from a more diverse perspective.
5.4. Comparison with Other Fusion Approaches
Table 6 presents the comparison between TransM3E and existing multi-modality fusion approaches, i.e., LateFuionSum, LateFusionConcat, and LateFusionMul. We compare our proposed method to these conventional multi-modal fusion approaches to illustrate that the performance improvement of our approach is not solely delivered by using the feature fusion between the skeleton modality and the RGB video modality. Compared with the best-performing baseline LateFusionConcat, our approach achieves a better performance.
@@ 6. Conclusion
In this paper, we propose the new task of video-based activated muscle group estimation in the wild. We contribute the first large-scale video-based AMGE dataset with in-the-wild videos and establish the MuscleMap benchmark using statistical baselines and existing video- and skeleton-based methods. Considering AMGE generalizability, we propose TransM3E with multi-classification token distillation and fusion in a cross-modality manner to enhance generalization to new activity types. TransM3E sets the state-of-the-art on the MuscleMap benchmark. Future works will explore missing-modality AMGE and leverage shared encoder.
Acknowledgement
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation) with a start date of April 1, 2024. This work was also partially supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Wuerttemberg and by the Federal Ministry of Education and Research. The authors also acknowledge support by the state of Baden-Wuerttemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This project is also supported by the National Key RD Program under Grant 2022YFB4701400.
@@ Appendix A Society Impact and Limitations
In our work, a new dataset targeting the AMGE is collected based on YouTube videos, termed MuscleMap135. We build up the MuscleMap benchmark for the AMGE by using statistic baselines and existing video-based approaches including both video-based and skeleton-based methods, while the three aforementioned datasets are all considered. Through the experiments, we find that the generalizability targeting AMGE on new activities is not satisfied for the existing activity recognition approaches. In order to tackle this issue, we propose a new cross-modality knowledge distillation approach named TransM3E while using MViTv2-S (Li et al., 2022) as its basic backbone. The proposed approach alleviates the generalization problem to a certain degree, however, there is still a large space for further improvement and future research. The AMGE performance gap between the known activities and new activities illustrates that our model has the potential to give offensive predictions, misclassification, and biased content which may cause false predictions resulting in a negative social impact. The dataset and code will be released publicly.
@@ Appendix B More details of the Dataset
The muscle regions where the number of sources is bigger than the threshold are chosen as activated muscle regions. We can see that no obvious deviation could be found in the AMGE annotation. We annotate the commonly leveraged human body muscles in daily life into $20$ muscle regions according to the suggestion of the experts, i.e., neck and head region, chest region, shoulder region, biceps region, triceps region, forearms region, upper back region, latissimus region, obliques region, upper abdominis region, lower abdominis region, lower back region, hamstring region, quadriceps region, calves region, inner thigh region, outer thigh region, gluteus region, feet ankles region, and wrists region. We rearrange occipitofrontalis, temporoparientalis, levator labii superioris, masticatorii, sternocleidomastoideus as neck and head muscle region; pectoralis major as chest region; deltoideus as shoulder region; biceps brachii as biceps region; triceps brachii as triceps region; flexor carpi radialis, palmaris longus, abductor pollicis longus as forearm region; trapezius as upper back region; latissimus dorsi as latissimus region; external oblique, serratus anterior as obliques region; rectus abdominis, quadratus lumborum as upper abdominis region; transversus abdominis, pyramidalis as lower abdominis region; erector spinae as lower back region; biceps femoris, semimembranosus, semitendinosus as hamstring region; rectus femoris, vastus medialis as quadriceps region; gastrocnemius, soleus as calves region; adductor longus, sartorius, gracilis as inner thigh region; iliotibial tract as outer thigh region; gluteus maximus as gluteus region; peroneus longus and brevis, extensor digitorum longus, flexor hallucis longus, flexor digitorum longus, peroneus tertius, tibialis posterior as feet ankles region; extensor pollicis, 1st dorsal interosseous, pronator quadratus as wrists region.
@@ Appendix E More Details of the MCTKD
Since we introduced the ablation regarding MCTKD in our main paper with experimental results, only more details regarding the KD format and position will be introduced in this section. In order to make it clearer for understanding, we illustrate more details regarding the KD/MCTKD position in Figure 6 to give a detailed clarification. For the MCTKD-related approaches, we use the MCTKD as depicted by (d), where the KD is executed between the knowledge receiver MCTs of the main modality and the sender MCTs of the auxiliary modality. For all the other basic KD-based approaches, we use the format as depicted by (c), where the KD is executed between the MCTs of the main modality and the MCTs of the auxiliary modality, regarded as conventional KD. All the experiments are executed with MCTs while without MCTF aggregation. We simply average the MCTs for all the experiments in this ablation. Regarding the sparse format as depicted in (a), the knowledge of the auxiliary modality is only transferred after the size reduction of the pooling layer denoted as DownSampling (DS) in Figure 6 and after the final layer. Only SparseMCTKD and DenseMCTKD are depicted since the SparseKD and DenseKD use the same position settings. SparseKD/MCTKD aims at reducing the KD/MCTKD calculation by selecting the most important intermediate layers to transfer the knowledge. After each pooling layer that has size reduction, the informative cues will be highlighted, which makes the corresponding changes of the tokens from auxiliary modality necessary to be integrated through KD/MCTKD. We choose the position after the pooling with size reduction to do the KD/MCTKD on the intermediate layer. DenseKD/MCTKD is designed to transfer the knowledge directly after each transformer block to leverage the knowledge from the other modality thoroughly. We make use of both KD positions to conduct a comparison and select the most appropriate method to build the MCTKD in our final model.